OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12. y = -12 - 4/3 x y = -4 - 4/3 x y = -12 + 4x y = -4 + 4x

OpenStudy (anonymous):

All you have to do is get y alone. First we can take a 4x off of each side. so minus a 4x from the left which will take that out and then to make it equal you must also minus a 4x from that side as well which gives us 3y = -12 -4x. Now to get that y alone by itself you have to divide each side by the 3 that's tagging along with the y so you get y = (-4x - 12)/3. Now divide each number on the right by 6 and you get y= -4/3x - 4 or, y = -4 - 4/3x, the second choice.

hero (hero):

4x + 3y = -12 3y = -12 - 4x y = (-12 - 4x)/3 y = -4 - 4x/3